OpenStudy (anonymous):

Simplify: square root of 54

OpenStudy (anonymous):

54 = 9*6

OpenStudy (anonymous):

square root those factors if you can.

OpenStudy (saifoo.khan):

\[\Large \sqrt{54} \to \sqrt{9 \times 6} \to 3 \sqrt6\]

OpenStudy (anonymous):

like the master^^^

OpenStudy (anonymous):

I don't get it I know 9*6 is 54 but how do you get the answer??

OpenStudy (saifoo.khan):

Ok, so first of all factor 54. you factored and got 9*6 now that's inside the square root. following on.. sqrt of 9 = 3 so, \[\sqrt{54} \to \sqrt{9 \times 6} \to \sqrt{3 \times 3 \times 6} \to 3 \sqrt6\]

OpenStudy (anonymous):

ok i think i get it thanks

OpenStudy (saifoo.khan):

Cool.